Hello, kind tutor!
Having a work problem concerning dilutions, and therefore ratios.
We took 30 microliters of a stock, and added it to 270 microliters of saline to make 300 microliters of a 1:10 dilution. We then took this 300 microliters and added 100 microliters of another solution to that to form 400 microliters total.
2 questions regarding this: What would the second dilution (300 microliters + 100 microliters) be considered? 3:4?
And, what would the final dilution be?
My boss and another coworker said it would end up being 1:13.3, (roughly 1:13), but I could not understand how they arrived at that number.
Thank you very much for your kind assistance,
Lauren

Answer by stanbon(75887)

Having a work problem concerning dilutions, and therefore ratios.
-----------------------------
We took 30 microliters of a stock, and added it to 270 microliters of saline to make 300 microliters of a 1:10 dilution.
We then took this 300 microliters and added 100 microliters of another solution (with no active ingrediant) to that to form 400 microliters total.

2 questions regarding this: What would the second dilution (300 microliters + 100 microliters) be considered? 3:4?
And, what would the final dilution be?
-----------------------
Equation:
active + active = active
0.10*300 + 0*100 = x(300+100)
30 + 0 = 400x
x = 30/400
x = 3/40
The ratio of active ingrediant to non-active in the final solution is 3:40
or 1:13 1/3
------
Congratualate your co-workers
=================
Cheers,
Stan H.